# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v2

@@ 3.5 Metastable states and chemical reactions
In previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@ 4 Fundamental constants and entropy increase
In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the entropy. Long-lived metastable states were important in this discussion because at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These long-lived metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and long-lived metastable phases emerging, with the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism (see Ref. Trachenko (2023d) for the review of related earlier discussions) where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability (fluctuations) of FPCs, analogous to genetic mutation in biological evolution, is currently unknown, although several mechanisms were discussed (see, e.g., Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013)).
@@
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.

### v3

@@ 3.5 Metastable states and chemical reactions
In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
@@ 4 Fundamental constants and entropy increase
In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy. We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
We recall that the discussion in Section 3 included long-lived metastable states. We saw that at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g. Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013) and Ref. Trachenko (2023d) for review).
@@
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space and entropy or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (similarly to biological evolution, this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
An example related to this discussion and involving constraints is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), breaking many constraints required for key processes of structure formation and disrupting these processes Barrow (2003); Adams (2019). This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs). We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.